OpenStudy (anonymous):

It takes Pam 45 minutes to drive to work and 60 minutes to drive home from work. Write the ratio of the time Pam spends driving home from work to the time she spends driving to work in three different ways.

OpenStudy (anonymous):

no

OpenStudy (hannahc234):

60/45 = 4/3 = 12/9